OpenStudy (math2400):

QUANTITATIVE If one species (2n = 16) crosses with another species (2n = 22), producing an allopolyploid offspring, what is the ploidy of the offspring?

OpenStudy (math2400):

@thomaster could you help me with this one?

OpenStudy (math2400):

answer choices are as follows: 2n = 8 + 11 = 19 2n = 22 4n = 32 + 44 = 76 2n = 16